# The coarse Baum-Connes conjecture with filtered coefficients and product metric spaces

### v1

@@ 5. Applications
In this section, we state that three ways mentioned in Introduction to the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients, thus, their product metric spaces satisfy the conjecture with filtered coefficients by Theorem 4.19. The reason is that all morphisms appeared in the proofs of those three ways are the identity map on filtered coefficients. The results in this section recover Fukaya and Oguni’s result in [5] and also contain some new examples satisfying the coarse Baum-Connes conjecture.
@@ 5.1. Metric spaces with finite asymptotic dimension
Remark 5.2.
Every polycyclic group (which is solvable) equipped with a proper metric has finite asymptotic dimension (cf. [13, Chapter 2]).
In [21], Yu invented the quantitative $K$-theory to prove that the coarse Baum-Connes conjecture holds for metric spaces with finite asymptotic dimension.
Let $X$ be a proper metric space and $A$ be a filtered $C^{\ast}$-algebra, then the Roe algebra $C_{f}^{\ast}(X,A)$ is also a filtered $C^{\ast}$-algebra equipped with a filtration defined by
@@
for $s\geq 0$ (please compare it with Lemma 3.9). Then by using the quantitative $K$-theory to $C_{f}^{\ast}(X,A)$ with the above filtration, we can obtain the following result.
@@ 5.2. Metric spaces with coarse embeddability into Hilbert space and CE-by-CE extensions
A proper metric space $X$ is called to be coarsely embedded into Hilbert space $H$, if there exists a map $f:X\rightarrow H$ and two non-deceasing functions $\rho_{\pm}:[0,\infty)\rightarrow[0,\infty)$ with $\lim_{t\rightarrow\infty}\rho_{-}(t)=\infty$ such that
@@
In [22], Yu used the Dirac-dual-Dirac method to prove that the coarse Baum-Connes conjecture holds for any metric space which admits a coarse embedding into Hilbert space (see [18, Chapter 12] for a simplified proof). We can take tensor products of the Dirac map and the dual-Dirac map with the identity map on filtered coefficients in the Dirac-dual-Dirac method, then we obtain the following result.
@@
The following notion was introduced by Deng, Wang and Yu in [2] in order to consider the coarse Baum-Connes conjecture for relative expanders. Recall that the coarse union of a sequence of metric spaces $(X_{i})_{i\in\mathbb{N}}$ is a disjoint union $\sqcup_{i\in\mathbb{N}}X_{i}$ equipped with a metric $d$ that restricts to the original metric on each $X_{i}$ and satisfies $d(X_{n},(\sqcup_{\mathbb{N}}X_{i})\setminus X_{n})\rightarrow\infty$ as $n\rightarrow\infty$.
@@
Using the same arguments (relies on the Dirac-dual -Dirac method) in [2], we can obtain the following lemma.
@@
Recently, Deng and Guo in [1] introduced a twisted coarse Baum-Connes conjecture and showed that the coarse Baum-Connes conjecture holds for the coarse union of $(\Gamma_{m})_{m\in\mathbb{N}}$ if coarse unions of $(N_{m})_{m\in\mathbb{N}}$ and $(Q_{m})_{m\in\mathbb{N}}$ satisfy the twisted coarse Baum-Connes conjecture with coefficients which generalizes the main result in [2]. Actually, we could obtain the similar result for the coarse Baum-Connes conjecture with filtered coefficients by considering the uniform version of the conjecture with filtered coefficients.
@@ 5.3. Open cones, coarse homotopy and relatively hyperbolic groups
Let $(M,d_{M})$ be a compact metric space with diameter at most two. The open cone over $M$, denoted by $\mathcal{O}M$, is a metric space defined to be the quotient space $(\mathbb{R}_{\geq 0}\times M)/(\{0\}\times M)$ with the following metric
@@
[6, Definition 2.1] Let $f,g:X\rightarrow Y$ be two coarse maps between proper metric spaces. $f$ and $g$ are called to be coarsely homotopic, if there exists a metric subspace $Z=\{(x,t):0\leq t\leq t_{x}\}$ of $X\times\mathbb{R}$ and a coarse map $h:Z\rightarrow Y$, such that
@@
The map $f$ is called a coarse homotopy equivalence map if there exists a coarse map $f^{\prime}:Y\rightarrow X$ such that $f^{\prime}f$ and $ff^{\prime}$ are coarsely homotopic to the identities $id_{X}$ and $id_{Y}$, respectively. Call $X$ and $Y$ be coarsely homotopy equivalent if there exists a coarse homotopy equivalence map from $X$ to $Y$.
Every open cone satisfies the coarse Baum-Connes conjecture and the proof relied on an Eilenberg swindle can be applied to the conjecture with filtered coefficients in any filtered $C^{\ast}$-algebras (cf. [24], [8]). Moreover, the conjecture with filtered coefficients is a coarse homotopy invariant (the proof relies on the Mayer-Vietoris six-term exact sequence, cf. [24], [9, Chapter 12]). Thus we have the following theorem.
@@
    Busemann non-positively curved spaces, for example CAT(0)-spaces;
@@
In [4], Fukaya and Oguni proved that if a group is hyperbolic relative to a finite family of infinite subgroups (cf. [4, Definition 2.4]), then this group satisfies the coarse Baum-Connes conjecture provided each subgroup belonging to the family satisfies the conjecture and admits a finite universal space for proper actions. By a similar argument, we can prove the following lemma.
@@
In [5], Fukaya and Oguni showed that the coarse Baum-Connes conjecture holds for products of hyperbolic groups, CAT(00)-groups, polyclic groups and certain relatively hyperbolic groups by consider the coronae of products. But it is generally not known whether the coarse Baum-Connes conjecture is true or not for products when every component satisfies the conjecture.
To sum up this section, by Theorem 4.19, we can obtain that the product spaces of metric spaces appeared in this section satisfy the coarse Baum-Connes conjecture with filtered coefficients, in particular, satisfy the original conjecture. That not only recovers Fukaya and Oguni’s result in [5, Theorem 1] by combining Examples 5.12 with Proposition 5.11 and Proposition 5.13, but also gives some new examples of products satisfying the coarse Baum-Connes conjecture.

### v2

@@ 5. Applications
In this section, we illustrate that three ways mentioned in the introduction to the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients. The reason is that all morphisms appeared in the proofs of those three ways are the identity map on filtered coefficients. The results in this section recover Fukaya and Oguni’s result in [6, Theorem 1.1] and also provide some new examples of product metric spaces satisfying the coarse Baum-Connes conjecture.
@@ 5.1. Metric spaces with finite asymptotic dimension
Example 5.2.
Every polycyclic group (which is solvable) equipped with a proper metric has finite asymptotic dimension (cf. [14, Chapter 2]).
In [24], Yu introduced the quantitative $K$-theory to prove that the coarse Baum-Connes conjecture holds for metric spaces with finite asymptotic dimension.
Let $X$ be a proper metric space and $A$ be a filtered $C^{\ast}$-algebra. Then the Roe algebra $C_{f}^{\ast}(X,A)$ is also a filtered $C^{\ast}$-algebra equipped with the filtration defined by
@@
for each $s\geq 0$ (please compare it with Lemma 3.9). Then by using the quantitative $K$-theory to $C_{f}^{\ast}(X,A)$ with the above filtration, we can obtain the following result.
@@ 5.2. Metric spaces with coarse embeddability into Hilbert space and CE-by-CE extensions
A proper metric space $X$ is called to be coarsely embedded into Hilbert space $H$, if there exists a map $f:X\rightarrow H$ and two non-decreasing functions $\rho_{\pm}:[0,\infty)\rightarrow[0,\infty)$ with $\lim_{t\rightarrow\infty}\rho_{-}(t)=\infty$ such that
@@
Example 5.5.
The class of metric spaces which admit a coarse embedding into Hilbert space contains hyperbolic spaces, mapping class groups and linear groups equipped with the word length metric (cf. [14]).
In [25], Yu used the Dirac-dual-Dirac method to prove that the coarse Baum-Connes conjecture holds for any metric space which admits a coarse embedding into Hilbert space (see [21, Chapter 12] for a simplified proof). We can take tensor products of the Dirac map and the dual-Dirac map with the identity map on filtered coefficients in the Dirac-dual-Dirac method, then we obtain the following result.
@@
The following notion was introduced by Deng, Wang and Yu in [3] in order to consider the coarse Baum-Connes conjecture for some relative expanders. Recall that the coarse union of a sequence of metric spaces $(X_{i})_{i\in\mathbb{N}}$ is a disjoint union $\sqcup_{i\in\mathbb{N}}X_{i}$ equipped with a metric $d$ that restricts to the original metric on each $X_{i}$ and satisfies $d(X_{n},(\sqcup_{\mathbb{N}}X_{i})\setminus X_{n})\rightarrow\infty$ as $n\rightarrow\infty$.
@@
Example 5.8.
Some relative expanders provide examples of CE-by-CE extensions, such as the box spaces of groups $\mathbb{Z}^{2}\rtimes_{Q}\mathbb{F}_{3}$ and $\mathbb{Z}\wr_{Q}\mathbb{F}_{3}$, where $Q$ is the kernel of the quotient map from $SL(2,\mathbb{Z})$ to $SL(2,\mathbb{Z}_{2})$ (cf. [3]).
Using the same arguments (relies on the Dirac-dual-Dirac method) in [3], we can obtain the following proposition.
@@
Recently, Deng and Guo in [2] introduced a twisted coarse Baum-Connes conjecture and they showed that the coarse Baum-Connes conjecture holds for the coarse union of $(\Gamma_{m})_{m\in\mathbb{N}}$ if coarse unions of $(N_{m})_{m\in\mathbb{N}}$ and $(Q_{m})_{m\in\mathbb{N}}$ satisfy the twisted coarse Baum-Connes conjecture with coefficients which generalizes the main result in [3]. Actually, we can obtain the similar result for the coarse Baum-Connes conjecture with filtered coefficients.
@@ 5.3. Open cones, coarse homotopy and relatively hyperbolic groups
Let $(M,d_{M})$ be a compact metric space with diameter at most two. The open cone over $M$, denoted by $\mathcal{O}M$, is defined to be the quotient space $(\mathbb{R}_{\geq 0}\times M)/(\{0\}\times M)$ with the following metric
@@
[7, Definition 2.1] Let $f,g:X\rightarrow Y$ be two coarse maps between proper metric spaces. $f$ and $g$ are called to be coarsely homotopic, if there exists a metric subspace $Z=\{(x,t):0\leq t\leq t_{x}\}$ of $X\times\mathbb{R}$ and a coarse map $h:Z\rightarrow Y$, such that
@@
The map $f$ is called a coarse homotopy equivalence map if there exists a coarse map $f^{\prime}:Y\rightarrow X$ such that $f^{\prime}f$ and $ff^{\prime}$ are coarsely homotopic to $id_{X}$ and $id_{Y}$, respectively. Call $X$ and $Y$ be coarsely homotopy equivalent if there exists a coarse homotopy equivalence map from $X$ to $Y$.
Every open cone satisfies the coarse Baum-Connes conjecture by an argument of the Eilenberg swindle. Its proof can be applied to the conjecture with filtered coefficients in any filtered $C^{\ast}$-algebras (cf. [27], [9]). Moreover, the conjecture with filtered coefficients is a coarse homotopy invariant (the proof relies on the Mayer-Vietoris six-term exact sequence, cf. [27], [10, Chapter 12]). Thus we have the following proposition.
@@
    Busemann non-positively curved spaces, for example, CAT(00)-spaces;
@@
Remark 5.15.
It is unknown whether every CAT(00)-space or Helly group admits a coarse embedding into Hilbert space.
In [5], Fukaya and Oguni proved that if a group is hyperbolic relative to a finite family of infinite subgroups (cf. [5, Definition 2.4]), then this group satisfies the coarse Baum-Connes conjecture provided each subgroup belonging to the family satisfies the conjecture and admits a finite universal space for proper actions. By a similar argument, we can prove the following proposition.
@@
In [6], Fukaya and Oguni showed that the coarse Baum-Connes conjecture holds for products of hyperbolic groups, CAT(00)-groups, polyclic groups and certain relatively hyperbolic groups by considering the coronae of products. But it is generally unknown whether the coarse Baum-Connes conjecture is true or not for products when their factors satisfy the conjecture.
To sum up this section, by Theorem 4.21, we can obtain that the products of metric spaces appeared in this section satisfy the coarse Baum-Connes conjecture with filtered coefficients. In particular, these products satisfy the coarse Baum-Connes conjecture. That not only recovers Fukaya and Oguni’s result in [6, Theorem 1] by combining Examples 5.14 with Proposition 5.13 and Proposition 5.16, but also gives some new examples of products satisfying the coarse Baum-Connes conjecture, such as the products of CAT(00)-groups or Helly groups with mapping class groups or linear groups equipped with the word length metric.